If 4x^4+9x^4=64 then the maximum value of x^2+y^2 is 

Grade:12

4 Answers

Akash Kumar Dutta
11 years ago

i think the ques. is 4x^4 + 9x^4=64...........................

Akash Kumar Dutta
11 years ago

from the eq. finding the value of x^2 and putting it in x^2 + y^2.we get 2nd eq.

differentiating that and putting it=0 to get the maximum value we get y^2=16/3.(13)^1/2

x^2= 12/(13)^1/2

hence. x^2 + y^2=52/3.(13)^1/2

                             =4.(root 13)/3     (ANS)

Rishabh Maggirwar
11 years ago

The value of x is 1.1, but what is the variable ''y''?

Abhishekh kumar sharma
11 years ago

the maximum value is 4(13)^2/3

substitute the value of eithe variable nad solve the equation x^2+y^2 and the get vale for which the equation is maximumthen put he value